# Jacobian Scopes: token-level causal attributions in LLMs

### v1

@@ 4 Discussion and conclusion
As summarized in Table˜1, the Jacobian Scopes range from logit-specific to fully distribution-aware, balancing clarity and explanatory scope with computational cost. Through various applications, we demonstrated the advantages of each formulation.
Semantic Scope attributes input tokens to a specific output logit or word choice. Fisher Scope attributes input tokens to the full predictive distribution, and is well suited for settings with multiple plausible outcomes, such as when translating, or where global uncertainty is of interest. Temperature Scope provides a computationally efficient alternative to Fisher Scope by attributing changes in distributional sharpness, making it suitable for long-context regimes where computing the full Jacobian is prohibitive. Section˜A.4 compares the attribution clarity of Fisher and Temperature Scopes on the same translation tasks.
More broadly, we hope to draw attention to the utility of Jacobian matrices in mechanistic interpretation of LLMs. A promising direction for future work is to develop additional Jacobian Scopes by exploring alternative projections and studying the spectral structure of the Jacobian and Fisher matrices.
@@
Linearized causality. Jacobian Scopes extract first-order input-output relationships via automatic differentiation, yielding local causal attributions in the linear neighborhood of the input. As such, they are closer in spirit to gradient-based methods such as SmoothGrad Smilkov et al. (2017) and Integrated Gradients Sundararajan et al. (2017); Sanyal and Ren (2021) than to explicitly interventional approaches like activation patching Heimersheim and Nanda (2024), and should not be interpreted as establishing global causal circuits Ameisen et al. (2025).
Architecture-blindness. Jacobian Scopes characterize input–output sensitivity without reference to a model’s internal transformer architecture. This makes the analysis more parsimonious, but also restricts the mechanistic insights Clark et al. (2019). As a result, Jacobian Scope attributions should be interpreted with awareness of potential circuit-level eccentricities at play Ameisen et al. (2025); Lieberum et al. (2024). For instance, Temperature Scope attributions in Figs.˜5 and 4 show elevated influence scores for early context tokens, an effect we attribute to the attention sink phenomenon Xiao et al. (2024). Section˜A.5 discusses further complications due to attention sink.
Dependence on back-propagation. Like most gradient-based attribution methods, Jacobian Scopes rely on back-propagation, which is more computationally intensive than forward-only approaches such as those based on pretrained auto-encoders Lieberum et al. (2024) or attention visualization Clark et al. (2019). For example, the Fisher Scope attribution for translating the French sentence in Fig.˜3 requires 153 seconds of CPU time, whereas its Temperature Scope counterpart in Fig.˜13 takes 1.4 seconds; by comparison, the forward pass alone takes merely 0.9 seconds. Therefore, large-scale analyses over extensive corpora are challenging without substantial GPU resources.
@@ Appendix A Appendix
A.1 Additional experiments on longer contexts
@@
To demonstrate the prowess of Jacobian Scopes on analyzing long-range interaction across numerous tokens, we provide additional examples on a larger model, LLaMA 3.2 3B.
@@
Temperature Scope: translating Dante. Accurate translation of literary text challenges a language model’s ability to reason coherently across many tokens. This makes translation ideal for probing whether Jacobian Scopes can faithfully reveal how meaning is analyzed across context. We prompt the LLM to perform translations by providing it with the following structure: ‘‘name of original language’’: [original text]. ‘‘name of target language’’: Fig.˜6 shows an example of LLaMA translating Dante’s Inferno Alighieri et al. (2002).
@@
Figure 6: Temperature Scope visualizes LLaMA-3.2 3B’s strategy for translating Dante. Left: Magnitude of influence from each word (blue) on predicted distribution at target position (red). Right: Probability distribution of top 7 words predicted at target position. (a) LLaMA assigns high probability to the token “world”, the unambiguous literal translation for the source world “mondo” in Italian. (b) LLaMA simultaneously adopts multiple strategies for translating the entire phrase “priegoti ch’a la mente altrui mi rechi”. Literally: “I pray thee to bring me to another’s mind.”
@@
As Fig.˜6 (b) shows, due to grammatical differences, a phrase in Italian may have multiple equally valid translations. When faced with such ambiguity, the LLM’s predicted next-word distribution becomes highly uncertain, with each one of the top-k tokens hinting at a different way to translate the original phrase. Therefore, we use Temperature Scope to visualize the part of the inputs that informed the entire predicted distribution, instead of just a certain word’s logit. The attribution graphs reveal two non-exclusive strategies: (a) when the source text has precise word-level counterparts to the target language, the LLM opts for word-to-word translation, with each translated word exactly attributed to its counterpart, and (b) if word-to-word translation is not possible due to mismatching syntactical structures, the LLM works on the level of phrases or sentences, with each predicted position attending to multiple tokens in the source language.
@@
A.2 Dynamical systems: definitions and additional experiments
@@
Figure 9: Temperature Scope attribution graph of LLaMA-3.2 extrapolating a partially observed Lorenz system, with different initial condition than Fig.˜4.
Figs.˜4 and 9 illustrate the model’s behavior under two different initial conditions. In Fig.˜4, the cutoff occurs at a peak in the $x$ value, while in Fig.˜9, the cutoff coincides with the $x$ value transitioning between attractor lobes. In both scenarios, the LLM attends to similar patterns surrounding the cutoff region.
Lorenz system with drift. To verify whether LLMs overlook the part of in-context data that are statistical outliers, we add a uniform drift term in the same Lorenz system shown in Fig.˜4. As such a system evolves, the later part of the contexts will eventually wander out of distribution compared to the earlier part. The resulting attribution pattern in Fig.˜10 indeed shows diminished influence score at the earlier regions in the context window.
@@
LLM attends to spurious patterns in Brownian motion. Notably, while the overall influence trend increases with context length — as seen in Fig.˜11 — there are many abrupt changes and irregularities in the influence score. These anomalies can be explained by random fluctuations: occasionally, patterns or trends may reappear in Brownian motion purely by statistical chance. For instance, in panel (b) of Section˜3.3, there is a sudden drop in $x$-values around token positions 100, which is“echoed" by a similar drop around the cutoff points at 360. Much like an amateur stock trader who perceives random fluctuations in stock movements as meaningful signals, the LLM incorrectly interprets these coincidental recurrences as informative patterns.
@@
A.3 Semantic vs. Temperature Scopes
Figure 12: Left: (a) Temperature Scope and (b) Semantic Scope attributions for the same Lorenz system. Right: The predicted probability mass over the top 20 candidate tokens at the target position, ordered by their numerical values. The ground-truth token (shown in red) is used as the target for Semantic Scope.
Both Semantic and Temperature Scopes define an influence score via a subspace projection of the input–output Jacobian, as formalized in Eq.˜1:
$\mathrm{Influence}_{t}:=\left\lVert\bm{v}^{\intercal}\bm{J}_{t}\right\rVert_{2},$
where the projection direction is chosen as $\bm{v}=\hat{\bm{y}}$ for Temperature Scope and $\bm{v}=\bm{w}_{\mathrm{target}}$ for Semantic Scope. The distinction therefore lies entirely in the choice of explanandum, i.e. the quantity to be explained: Semantic Scope attributes influence with respect to a single target logit, whereas Temperature Scope attributes influence with respect to the predictive distribution as a whole.
@@
Naively applying Semantic Scope with the ground-truth token (e.g. “84”) as the target yields misleading attributions: the resulting influence map highlights only the inputs that increase the logit of “84”, while ignoring the remaining tokens that collectively constitute the bell-shaped predictive distribution. As a result, the attribution fails to capture the model’s reasoning at the distributional level. Temperature Scope avoids this failure mode by attributing influence with respect to the predicted distribution as a whole, rather than any single outcome. This makes it better aligned with forecasting settings in which uncertainty and variability are intrinsic to the task.
A.4 Fisher vs. Temperature Scopes
Both Fisher and Temperature Scopes attribute input tokens to the full predictive distribution. However, Fisher Scope is substantially more expensive in terms of compute, while also producing cleaner and more interpretable attribution patterns. Fig. 13 illustrates this cost – clarity trade-off by comparing the two Scopes on the same translation task with LLaMA-3.2 1B.
Figure 13: Fisher Scope (top) versus Temperature Scope (bottom) attributions for the same set of translation tasks. For the French example shown, Fisher Scope requires 153 seconds to evaluate, compared to 1.4 seconds for Temperature Scope.
As shown in Fig.˜13, Fisher Scope yields highly precise attribution patterns that assign each translated word to its corresponding (sub)word in the source language; for example, researchers $\rightarrow$ ricercatori, attribution $\rightarrow$ Zuschreibung, and Scope $\rightarrow$ lentille. In contrast, Temperature Scope assigns influence more diffusely, highlighting these key source words alongside additional contextual tokens that collectively shape the predictive distribution.
A.5 Path-Integrated Semantic Scope

### v2

@@ 5 Discussion and conclusion
As summarized in Table˜1, the three Jacobian Scopes span a spectrum from logit-specific (Semantic) to fully distribution-aware (Fisher, Temperature), each targeting a distinct facet of the model’s output: the logit of a specific token, the shape of the full predictive distribution, and the sharpness of that distribution, respectively.
More broadly, we put forward vector-Jacobian projection as a unifying framework for gradient-based attribution in LLMs. A profitable direction for future work is to derive new explananda by studying the spectral structure of the Jacobian $\bm{J}_{t}$ and the Fisher information matrix $\bm{F}$.
@@
Linearized causality. Jacobian Scopes extract first-order input-output relationships via automatic differentiation, yielding local causal attributions in the linear neighborhood of the input. As such, they are closer in spirit to gradient-based methods such as SmoothGrad Smilkov et al. (2017) and Integrated Gradients Sundararajan et al. (2017); Sanyal and Ren (2021) than to explicitly interventional approaches like activation patching Heimersheim and Nanda (2024); Ghandeharioun et al. (2024) or circuit tracing Ameisen et al. (2025).
Architecture-blindness. Jacobian Scopes characterize input–output sensitivity without reference to a model’s internal transformer architecture. This makes the analysis more parsimonious, but also restricts the mechanistic insights Clark et al. (2019). As a result, Jacobian Scope attributions should be interpreted with awareness of potential circuit-level eccentricities at play Ameisen et al. (2025); Lieberum et al. (2024). For instance, Temperature Scope attributions in Fig.˜4 (b) show elevated influence scores for early context tokens, an effect we attribute to the attention sink phenomenon Xiao et al. (2024). Section˜A.7 discusses further complications due to attention sink.
Dependence on back-propagation. Like most gradient-based attribution methods, Jacobian Scopes rely on back-propagation, which is in theory more computationally intensive than forward-only approaches such as those based on pretrained auto-encoders Lieberum et al. (2024) or attention visualization Clark et al. (2019).
In practice, however, the overhead is negligible: on a single NVIDIA RTX A4000, the backward pass for the example in Fig.˜3 takes only 0.027 seconds, compared to 0.069 seconds for the forward pass. This efficiency follows directly from the design of Jacobian Scopes: rather than backpropagating through all model parameters, each attribution requires only a single vector-Jacobian product with respect to the input embeddings, with parameter gradients disabled throughout.
@@ Appendix A Appendix
A.1 Additional information on Benchmarks
Dataset snapshots. Figure˜5 shows Fisher Scope attributions for 1 out of the 1000 passages of each dataset using LLaMA-3.2 3B.
[…]
Full quantitative results.
Table 3: AOPC (higher magnitude is better) for LLaMA-3.2 1B, averaged over ablation fractions $k\in\{5\%,10\%,20\%\}$ via the trapezoidal rule over 1000 test passages.
Method LAMBADA IWSLT2017
Random $0.26\pm 0.01$ $0.30\pm 0.01$
Integrated Grads $1.10\pm 0.01$ $0.93\pm 0.01$
Input $\times$ Grad $1.28\pm 0.01$ $1.04\pm 0.01$
Semantic Scope $1.30\pm 0.01$ $1.06\pm 0.01$
Temperature Scope $\mathbf{-1.32\pm 0.01}$ $\mathbf{-1.09\pm 0.01}$
Fisher Scope $\mathbf{-1.32\pm 0.01}$ $1.08\pm 0.01$
Table 4: AOPC (higher magnitude is better) for Qwen2.5-1.5B, averaged over ablation fractions $k\in\{5\%,10\%,20\%\}$ via the trapezoidal rule over 1000 test passages.
Method LAMBADA IWSLT2017
Random $0.35\pm 0.01$ $0.25\pm 0.01$
Integrated Grads $1.71\pm 0.01$ $1.35\pm 0.01$
Input $\times$ Grad $1.72\pm 0.01$ $1.35\pm 0.01$
Semantic Scope $1.68\pm 0.01$ $1.24\pm 0.01$
Temperature Scope $\mathbf{-1.86\pm 0.01}$ $\mathbf{-1.38\pm 0.01}$
Fisher Scope $1.77\pm 0.01$ $\mathbf{-1.38\pm 0.01}$
Table 5: AOPC (higher magnitude is better) for Qwen2.5-3B, averaged over ablation fractions $k\in\{5\%,10\%,20\%\}$ via the trapezoidal rule over 1000 test passages.
Method LAMBADA IWSLT2017
Random $0.29\pm 0.01$ $0.20\pm 0.01$
Integrated Grads $39\pm 0.01 1.1$$11\pm 0.01$
Input $\times$ Grad $39\pm 0.01 1.1$$11\pm 0.01$
Semantic Scope $1.34\pm 0.01$ $1.01\pm 0.02$
Temperature Scope $\mathbf{-1.55\pm 0.01}$ $\mathbf{-1.20\pm 0.01}$
Fisher Scope $41\pm 0.01 1.1$$15\pm 0.01$
Tables˜3, 4 and 5 report AOPC for LLaMA-3.2 1B, Qwen2.5-1.5B, and Qwen2.5-3B respectively, complementing the LLaMA-3.2 3B results in Table˜2.
Across all models and datasets, Jacobian Scopes consistently outperform random ablation and Integrated Gradients, and perform on par with or better than Input $\times$ Gradient.
On LLaMA models, Fisher Scope and Temperature Scope perform comparably, with Fisher Scope showing a slight edge on IWSLT2017 — consistent with its design for tasks with non-unique predictions.
On Qwen models, however, Temperature Scope emerges as the strongest performer across both datasets, suggesting that the relative strengths of the scopes exhibit mild model dependence. We leave a deeper investigation of this phenomenon to future work.
Datasets and preprocessing. LAMBADA Paperno et al. (2016) is a dataset of narrative passages specifically designed so that a human reader would need to process the full context in order to predict the final word — local context alone is insufficient. This property makes it ideal for evaluating token-level attribution methods, as a faithful attribution should identify long-range contextual dependencies rather than trivial local patterns. We retain only passages of at least 50 words, and remove any trailing punctuation from the final word.
[…]
Evaluation metric. We evaluate using the Area Over the Perturbation Curve (AOPC) Samek et al. (2016), which measures attribution faithfulness by progressively ablating the most influential tokens and recording the resulting drop in the ground-truth target token’s log-probability. Ablated tokens are replaced with zero vectors. Formally, for ablation fractions $k\in\{5\%,10\%,20\%\}$, the AOPC is computed using the trapezoidal rule
$\mathrm{AOPC}=\sum_{i}\frac{1}{2}(k_{i+1}-k_{i})\left(\Delta\log p_{k_{i}}+\Delta\log p_{k_{i+1}}\right),$ (7)
where $\Delta\log p_{k}$ denotes the drop in log-probability when the top-kk% most influential tokens are ablated. A more negative AOPC indicates that the ablated tokens were genuinely influential, reflecting higher attribution quality. The $\pm$ values in all tables report the standard error of the mean across the 1000 test passages.
Baselines. Input $\times$ Gradient Shrikumar et al. (2017) defines the influence score of token $\bm{x}_{t}$ as the $\ell_{2}$ norm of the element-wise product of the input embedding and its gradient with respect to the target logit:
$\mathrm{Influence}_{t}^{\mathrm{I\times G}}:=\left\|\bm{x}_{t}\odot\nabla_{\bm{x}_{t}}z_{\mathrm{target}}\right\|_{2}.$ (8)
Integrated Gradients Sundararajan et al. (2017), defined in Eq.˜11, integrates the gradient along a linear interpolation path from null (all zero) input to the full input. We refer the reader to Section˜A.7, which visualizes the path integration in detail and discusses why this approach introduces distortions in the LLM setting — most notably attention sink effects at small interpolation values — that likely contribute to its weaker AOPC performance.
A.2 Related work
Gradient-based attribution methods were first systematically developed for convolutional image classifiers. Simonyan et al. (2013) introduced class saliency maps by computing the gradient of the predicted class score with respect to the input pixels, identifying which image regions drive a given classification decision in a single backward pass. Building on this, Shrikumar et al. (2017) proposed DeepLIFT, which propagates neuron contribution scores by comparing each activation to a reference state; gradient $\times$ input—the elementwise product of the input embedding and its gradient with respect to the target logit—arises as a simpler special case and serves as a baseline in our evaluations.
A key challenge with plain gradient saliency is that the resulting maps are often noisy and fail to sharply localize the relevant input features. SmoothGrad (Smilkov et al., 2017) addresses this for vision models by averaging sensitivity maps over many copies of the input perturbed with Gaussian noise. Integrated Gradients (Sundararajan et al., 2017) instead eliminates path-dependence by integrating the gradient along a linear interpolation from a null baseline to the actual input, satisfying axiomatic desiderata such as Sensitivity and Implementation Invariance. Sanyal and Ren (2021) subsequently adapted this path integration to the discrete embedding space of language models.
All of these methods attribute a scalar model output—a single class logit or token log-likelihood—and are therefore conceptually analogous to Semantic Scope. Yet LLMs produce a full distribution over the vocabulary, and no prior methods, to our best knowledge, exploit this richer structure.
By contrast, Fisher Scope and Temperature Scope natively attribute the full predictive distribution at no additional computational cost: Fisher Scope uses the principal direction of the Fisher information matrix, which captures the single perturbation direction in hidden space that most changes the distribution as measured by KL divergence; Temperature Scope targets the effective inverse temperature, summarizing how sharply peaked the predictive distribution is.
This advantage is reflected in our benchmark results: across both LAMBADA and IWSLT2017, Fisher and Temperature Scopes consistently achieve the strongest AOPC scores among all methods compared.
A.3 Additional experiments on longer contexts
@@
To demonstrate the prowess of Jacobian Scopes on analyzing long-range interaction across numerous tokens, we provide additional examples on a larger model, LLaMA-3.2 3B.
@@
Fisher Scope: translating Dante. Accurate translation of literary text challenges a language model’s ability to reason coherently across many tokens. This makes translation ideal for probing whether Jacobian Scopes can faithfully reveal how meaning is analyzed across context. We prompt the LLM to perform translations by providing it with the following structure: ‘‘name of original language’’: [original text]. ‘‘name of target language’’: Fig.˜6 shows an example of LLaMA translating Dante’s Inferno Alighieri et al. (2002).
@@
Figure 6: Fisher Scope visualizes LLaMA-3.2 3B’s strategy for translating Dante. Left: Magnitude of influence from each word (blue) on predicted distribution at target position (red). Right: Probability distribution of top 7 words predicted at target position. (a) LLaMA assigns high probability to the token “world”, the unambiguous literal translation for the source world “mondo” in Italian. (b) LLaMA simultaneously adopts multiple strategies for translating the entire phrase “priegoti ch’a la mente altrui mi rechi”. Literally: “I pray thee to bring me to another’s mind.”
@@
As Fig.˜6 (b) shows, due to grammatical differences, a phrase in Italian may have multiple equally valid translations. When faced with such ambiguity, the LLM’s predicted next-word distribution becomes highly uncertain, with each one of the top-k tokens hinting at a different way to translate the original phrase. Fisher Scope is therefore ideal in such cases, as it highlights the inputs that informed not just a target logit, but instead the entire predicted distribution over all vocabulary.
The attribution graphs reveal two non-exclusive strategies: (a) when the source text has precise word-level counterparts to the target language, the LLM opts for word-to-word translation, with each translated word exactly attributed to its counterpart, and (b) if word-to-word translation is not possible due to mismatching syntactical structures, the LLM works on the level of phrases or sentences, with each predicted position attending to multiple tokens in the source language.
@@
A.4 Dynamical systems: definitions and additional experiments
@@
Figure 9: Temperature Scope attribution graph of LLaMA-3.2 extrapolating a partially observed Lorenz system, with different initial condition than Fig.˜4 (a).
Figs.˜4 and 9 illustrate the model’s behavior under two different initial conditions. In Fig.˜4 (a), the cutoff occurs at a peak in the $x$ value, while in Fig.˜9, the cutoff coincides with the $x$ value transitioning between attractor lobes. In both scenarios, the LLM attends to similar patterns surrounding the cutoff region.
Lorenz system with drift. To verify whether LLMs overlook the part of in-context data that are statistical outliers, we add a uniform drift term in the same Lorenz system shown in Fig.˜4 (a). As such a system evolves, the later part of the contexts will eventually wander out of distribution compared to the earlier part. The resulting attribution pattern in Fig.˜10 indeed shows diminished influence score at the earlier regions in the context window.
@@
LLM attends to spurious patterns in Brownian motion. Notably, while the overall influence trend increases with context length — as seen in Fig.˜11 — there are many abrupt changes and irregularities in the influence score. These anomalies can be explained by random fluctuations: occasionally, patterns or trends may reappear in Brownian motion purely by statistical chance. For instance, in panel (b) of Fig.˜4, there is a sudden drop in $x$-values around token positions 100, which is “echoed” by a similar drop around the cutoff points at 360. Much like an amateur stock trader who perceives random fluctuations in stock movements as meaningful signals, the LLM incorrectly interprets these coincidental recurrences as informative patterns.
@@
A.5 Temperature Scope is optimal for numerical data
@@
Figure 12: Left: (a) Temperature Scope (b) Semantic Scope and (c) Fisher Scope attributions for the same Lorenz system. Right: The predicted probability mass over the top 20 candidate tokens at the target position, ordered by their numerical values. The ground-truth token (shown in red) is used as the target for Semantic Scope.
Naively applying Semantic Scope with the ground-truth token (e.g. “84”) as the target yields misleading attributions: the resulting influence map highlights only the inputs that increase the logit of “84”, while ignoring the remaining tokens that collectively constitute the bell-shaped predictive distribution. As a result, the attribution fails to capture the model’s reasoning at the distributional level.
Fisher Scope also exhibits attenuated attribution signals at early context positions. A heuristic explanation is as follows. Fisher Scope quantifies how input perturbations affect the predictive distribution over the entire vocabulary, including both numerical and non-numerical tokens. However, the underlying dynamical system is defined only over numerical tokens. Early numerical inputs therefore inform the implicit transition rule learned by the LLM (Liu et al., 2024) and can significantly influence future numerical predictions. Yet these inputs carry little information about non-numerical tokens, whose probabilities tend to default to patterns determined by near-term context. Because Fisher Scope aggregates sensitivity across the entire vocabulary, this mismatch dilutes the attribution signal originating from early numerical tokens.
Temperature Scope, by contrast, is dominated by the most probable tokens, which in this setting are numerical tokens. It therefore focuses attribution on the portion of the distribution that is relevant to the forecasting task, effectively bypassing the out-of-support vocabulary. This makes Temperature Scope better aligned with forecasting scenarios in which uncertainty and variance are intrinsic to the prediction.
A.6 Effective temperature as variance
The bell-shaped predictive distribution over numerical tokens observed in Figs.˜12 and 4 offers a precise interpretation of $\beta_{\mathrm{eff}}$: under the approximation that the distribution is Gaussian, $\beta_{\mathrm{eff}}$ is proportional to the inverse predictive variance $\frac{1}{\sigma^{2}}$. We establish this connection via the following proposition.
Proposition 1.
Treat the numerical values $v$ of tokens as samples from a continuous distribution. Let $p(v)\propto e^{\,\beta\,\hat{z}(v)}$ be the predicted distribution, where $\hat{z}(v)$ is the logit assigned to the token with value $v$ and $\beta=\beta_{\mathrm{eff}}$. Then $p(v)$ is a Gaussian distribution if and only if $\hat{z}(v)$ is quadratic in $v$. Moreover, writing $\hat{z}(v)=-b(v-\mu)^{2}+c$ with $b>0$, the variance of $p(v)$ is
$\sigma^{2}=\frac{1}{2\beta b},$ (10)
so $\beta\propto\sigma^{-2}$, with $b$ a model-dependent constant.
Proof.
$(\Leftarrow)$ Suppose $\hat{z}(v)=-b(v-\mu)^{2}+c$. Then
$e^{\,\beta\,\hat{z}(v)}=e^{\beta c}\cdot e^{-\beta b(v-\mu)^{2}},$
which is a Gaussian kernel with mean $\mu$ and variance $\sigma^{2}=1/(2\beta b)$.
$(\Rightarrow)$ Suppose $p(v)$ is Gaussian with mean $\mu$ and variance $\sigma^{2}$, so that $e^{\,\beta\,\hat{z}(v)}\propto e^{-(v-\mu)^{2}/(2\sigma^{2})}$. Taking logarithms gives
$\hat{z}(v)=-\frac{1}{2\beta\sigma^{2}}(v-\mu)^{2}+\mathrm{const},$
which is quadratic with $b=1/(2\beta\sigma^{2})$, recovering $\sigma^{2}=1/(2\beta b)$. ∎
[…]
A.7 Path-Integrated Semantic Scope